OpenStudy (anonymous):

is the function f(x)=x^2+x+1 even odd or neither

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

Can you briefly explain why it is neither

OpenStudy (anonymous):

it is neither symmetric about origin nor about y-axis

OpenStudy (anonymous):

a function is even if: f(x) = f(-x) or in other words, a function is even if it has the same value for x or -x a function is odd if: f(x) + f(-x) = 0 or in other words, a function if the function evaluated at -x gives the opposite value of the function evaluated at x the function in your post is neither, because it doesn't fit the requirements.

OpenStudy (anonymous):

for instance: f(x)=x^2+x+1 f(1) = 1^2 + 1 + 1 = 3 f(-1) = 1 - 1 + 1 or 2 if it was even, f(1) would equal f(-1) [they would both be 3] if it was odd, f(1) would be the oppsite of f(-1) [one is 3, the other is -3] neither is the case here